OpenStudy (raffle_snaffle):

The number of people arriving for treatment at an emergency room can be modeled by a Poisson process with a rate parameter of six per hour. (a) What is the probability that exactly four arrivals occur during a particular hour? (Round your answer to three decimal places.) (b) What is the probability that at least four people arrive during a particular hour? (Round your answer to three decimal places.) (c) How many people do you expect to arrive during a 45-min period?

OpenStudy (raffle_snaffle):

My solution is attached

OpenStudy (raffle_snaffle):

@jim_thompson5910

OpenStudy (raffle_snaffle):

b and c are wrong. I got a right

jimthompson5910 (jim_thompson5910):

I'm getting 0.938 for b as well. Did you make sure to round to only 3 decimal places? I see that you've written out 5 decimal places

jimthompson5910 (jim_thompson5910):

oh wait, you were missing a term, you should have P(x = 3) in there as well

OpenStudy (raffle_snaffle):

When I put the answer in the box i put three = 0.938

OpenStudy (raffle_snaffle):

you are absolutely right.

jimthompson5910 (jim_thompson5910):

since P(x >= 4) = 1 - P(x < 4)

OpenStudy (raffle_snaffle):

0.849

jimthompson5910 (jim_thompson5910):

as for part c, maybe you just leave it as 4.5 or 4.500 and not round that to 5

jimthompson5910 (jim_thompson5910):

I'm getting 0.849 too

OpenStudy (raffle_snaffle):

okay what about C?

jimthompson5910 (jim_thompson5910):

for part c, maybe you just leave it as 4.5 or 4.500 and not round that to 5

OpenStudy (raffle_snaffle):

but you can't have 4.5 people lol

jimthompson5910 (jim_thompson5910):

it's an expected (ie average) number

jimthompson5910 (jim_thompson5910):

example day 1 you have 10 people day 2 you have 11 people on average, you have (10+11)/2 = 21/2 = 11.5 people

OpenStudy (raffle_snaffle):

Thanks that was all right

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (raffle_snaffle):

okay okay lol